# Control Barrier Functions via Minkowski Operations for Safe Navigation among Polytopic Sets

### v2

@@ Abstract
Safely navigating around obstacles while respecting the dynamics, control, and geometry of the underlying system is a key challenge in robotics. Control Barrier Functions (CBFs) generate safe control policies by considering system dynamics and geometry when calculating safe forward-invariant sets. Existing CBF-based methods often rely on conservative shape approximations, like spheres or ellipsoids, which have explicit and differentiable distance functions. In this paper, we propose an optimization-defined CBF that directly considers the exact Signed Distance Function (SDF) between a polytopic robot and polytopic obstacles. Inspired by the Gilbert-Johnson-Keerthi (GJK) algorithm, we formulate both (i) minimum distance and (ii) penetration depth between polytopic sets as convex optimization problems in the space of Minkowski difference operations (the $\mathcal{M}_{\mathcal{D}}$-space). Convenient geometric properties of the $\mathcal{M}_{\mathcal{D}}$-space enable the derivatives of implicit SDF between two polytopes to be computed via differentiable optimization. We demonstrate the proposed framework in three scenarios including pure translation, initialization inside an unsafe set, and multi-obstacle avoidance. These three scenarios highlight the generation of a non-conservative maneuver, a recovery after starting in collision, and the consideration of multiple obstacles via pairwise CBF constraint, respectively. A video of all simulation animations can be found at https://youtu.be/3Dh0gtDW8bE .
@@ I Introduction
Safely navigating an environment with obstacles is a fundamental challenge in robotics. Navigational solutions must respect the dynamics and control constraints of the vehicle and the geometric constraints of the vehicle and obstacles. Control Barrier Functions (CBFs) can help guarantee safety for control-affine systems by providing a forward-invariant set that respects dynamics, control, and geometry. CBFs can be combined with Control Lyapunov Functions (CLFs) to enable controllers that stabilize the system while respecting safety constraints and control limits [1, 2]. This combination yields a sequence of constrained optimization problems that, assuming quadratic costs, reduce to Quadratic Programs (QP). QPs are computationally efficient and well-suited to real-time applications [1]. Further advancements in safety-critical control with CBFs have introduced adaptive control and formulations for high relative degree constraints [3, 4, 5, 6, 7].
@@
CBFs are used for obstacle avoidance across a range of applications [8, 9, 10, 11, 12, 13, 14], yet most CBFs approximate robots and/or obstacles by smooth shapes such as hyperspheres [15] and ellipsoids [16, 17]. Hyperspheres and ellipsoids are popular because they enable differentiable closed-form expressions for robot-to-obstacle distance. On the other hand, such shapes are often overly conservative approximations of real-world objects that are, for example, more accurately approximated by polytopes.
@@
In this paper we present a CBF-based framework that considers polytopic robots and polytopic obstacles directly. Our method computes the exact signed distance function (SDF) between two polytopes by solving a convex optimization problem in the dual space that is defined by the Minkowski difference between the two polytopes (see Fig. 1). The resulting robot-to-obstacle SDF is differentiable almost everywhere.
@@
Our work builds on a foundation of prior work [18, 19, 20, 21, 22, 23] that considers continuous-time CBFs for collision avoidance between polytopes/polygons. The first study of CBFs [18] uses the Minimum Distance Function (MDF), solved by duality-based convex optimization, to construct a non-smooth CBF (NCBF). In [18, 19] the exact derivative is applied in the case of strongly convex sets (but not polytopes, which are weakly convex), while a conservative lower bound on the CBF derivative is used for polytopes. A CBF-based filtering strategy for manipulators [20] uses the SDF and a local linear approximation of the CBF gradient (approximation error considered to ensure safety). An optimization-free smooth approx-SDF CBF is proposed in [23], where conservatism is reduced via hyperparameter tuning. In [22] CBFs are constructed using differentiable optimization of the minimum scaling factor (another metric for collision detection), and polygons are approximated by strictly convex shapes. Polygons are similarly approximated by polynomial shapes using logistic regression in [21]. Our work differs from prior work [18, 19, 20, 21, 22, 23] in that we solve the SDF via convex optimization and compute its derivative using differentiable optimization without approximation by leveraging the configuration obstacle defined via the Minkowski difference. Table I (next page) provides a conceptual comparison with existing works.
The main contributions of this paper are as follows:
II Preliminaries
II-A Control Barrier Functions
Consider a nonlinear affine control system:
$\dot{\bm{x}}=f(\bm{x})+g(\bm{x})\bm{u},$ (1)
Definition 1 (Forward invariant set):
Let $C$ be the safe set, defined as the 0-superlevel set of a continuously differentiable function $h:\mathcal{X}\subset\mathbb{R}^{n}\rightarrow\mathbb{R}$ :
Definition 2 (CBF [2]):
Given a set $C$ as in (2), $h(\bm{x})$ is a control barrier function (CBF) for system (1) if there exists an extended class $\mathcal{K}_{\infty}$ function [^2] An extended class $\mathcal{K}_{\infty}$ function is a function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ that is strictly increasing and and $\alpha(0)=0$, see, e.g., [1]. $\alpha$ such that, for all $\bm{x}\in\mathcal{X}$,
where $L_{f},L_{g}$ denote the Lie derivatives along $f$ and $g$, respectively. It is assumed that $L_{g}h(\bm{x})\neq 0$ when $h(\bm{x})=0$.
Definition 3 (Control Lyapunov function [31]):
Theorem 1 (Safety Guarantee from [2]).
Given a CBF $h$ with the associated set $C$ from (2), any Lipschitz continuous controller $\bm{u}(t)\in\mathcal{U},t\geq t_{0}$ that satisfies (3) renders the set $C$ forward invariant for control system (1).
Many existing works [2, 32, 4], unify CBF and CLF within a quadratic programming (QP) framework. The CLF constraint is always relaxed using a slack variable, which is then minimized in the cost function to improve feasibility.
@@ II-B Collision Detection via Minkowski Operation
The workspace $\mathcal{W}$ of a robot is the physical (2D or 3D Euclidean) environment in which it operates, consisting of the space where the robot can move and the obstacles it must avoid. The configuration space ($\mathcal{C}$-space) represents all possible positions, orientations, etc. with each point defining a unique configuration. In contrast, the $\mathcal{M}_{\mathcal{D}}$-space is a dual of $\mathcal{W}$, and is a Euclidean space with the same (2D or 3D) dimensionality as $\mathcal{W}$. In $\mathcal{M}_{\mathcal{D}}$-space, the robot is shrunk to a point at the origin, and the obstacles are transformed to encode the robot’s geometry and configuration (see Fig. 1).
@@
where =$-\mathcal{R}=\{-\bm{p}\mid\bm{p}\in\mathcal{R}\}$ is the mirror image projection of $\mathcal{R}$ through the origin, in which $\bm{p}$ is a position vector of $\mathcal{R}$. Note that, in this paper, we refer to the Minkowski sum of $\mathcal{O}$ and $-\mathcal{R}$ as the Minkowski difference [^3] In fields like mathematics and image processing, the Minkowski difference has a different (alternative) definition, where $\mathcal{A}\ominus\mathcal{B}$ is the set that, when added to $\mathcal{B}$, recovers $\mathcal{A}$. of $\mathcal{O}$ and $\mathcal{R}$ to follow the convention in collision detection.
@@
While the Minkowski sum (5) and CO computation (6) are general operations that apply to any shape, it is worth emphasizing that Lem. 1 only holds for convex sets. When a non-convex set is well behaved (e.g., non-fractal) such that it can be decomposed into a union of convex sets, then a CO can be computed for each combination of a convex robot set with a convex obstacle set, and Lem. 1 applied iteratively across these interactions. In the following, without loss of generality, we assume both the robot and the obstacle are nonempty and convex polygons.
@@ II-C Signed distance for Collision Avoidance
The signed distance is positive when $\mathcal{R}$ and $\mathcal{O}$ are disjoint and negative when they intersect. Thus, collision avoidance can be enforced by ensuring $\text{sd}(\mathcal{R},\mathcal{O})>0$. The signed distance between two sets can equivalently be expressed as the signed distance between the origin $\bm{0}_{c}$ and their CO.
The closest related work [18] constructs CBFs based on implicit MDFs by formulating a convex optimization problem with dual variables. However, the resulting NCBF constraint is enforced using a lower bound of the CBF derivative via linear programming (LP). Furthermore, they consider only the minimum distance and do not account for penetration depth in cases where the system starts in collision. To address this issue, we propose a new framework for SDF-based CBFs that incorporates penetration depth, as detailed in Sec. IV.
@@ III Problem Formulation
The robot’s state is denoted $\bm{x}$. Both robot and obstacles are defined using the halfspace-representation (H-rep), $\mathcal{R}(\bm{x}):=\{y\in\mathbb{R}^{d}\mid A_{r}(\bm{x})y\leq b_{r}(\bm{x})\}$ and $\mathcal{O}_{i}:=\{y\in\mathbb{R}^{d}\mid A_{o_{i}}y\leq b_{o_{i}}\}$, with $A_{r}\in\mathbb{R}^{\ell_{r}\times d},\leavevmode\nobreak\ b_{r}\in\mathbb{R}^{\ell_{r}}$, $A_{o_{i}}\in\mathbb{R}^{\ell_{o_{i}}\times d},\leavevmode\nobreak\ b_{o_{i}}\in\mathbb{R}^{\ell_{o_{i}}}$, and $d\in\{2,3\}$ for 2D and 3D spaces. $\ell_{r}$ and $\ell_{o_{i}}$ denote the number of edges of the robot and the $i$-th obstacle, respectively. The corresponding COs are denoted as $\mathcal{O}_{i}^{\mathcal{C}}(\bm{x})=\{y\in\mathbb{R}^{d}\mid A^{\mathcal{C}}_{i}(\bm{x})y\leq b^{\mathcal{C}}_{i}(\bm{x})\}$, omitting the subscipt $\mathcal{R}$ for brevity, where $A_{i}^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}_{i}}\times d},\leavevmode\nobreak\ b_{i}^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}_{i}}}$, and $\ell_{\mathcal{C}_{i}}$ denotes the number of edges of the $i$-th CO.
@@
  2. 2)
    $A^{\mathcal{C}}_{i},b^{\mathcal{C}}_{i}$ are continuously differentiable almost everywhere, and the inequality constraints provide a minimal representation of the CO. For almost all $\bm{x}\in\mathcal{X}$, $\ell_{\mathcal{C}_{i}}=\ell_{r}+\ell_{o_{i}}$. Note that, if $\mathcal{R}$ and $\mathcal{O}_{i}$ have parallel edges, the corresponding constraints $A^{\mathcal{C}}_{i},\leavevmode\nobreak\ b^{\mathcal{C}}_{i}$ may lose one or more dimensions, depending on the number of parallel edges. However, this degenerated case occurs only on a measure zero set and, thus, a.e. does not affect the optimization result (as reported in [37]).
  3. 3)
    All obstacles are static, uniformly bounded and have nonempty interior and the information of obstacles is accessible to the robot.
@@
[…]
@@
This formulation extends to multiple obstacles by applying the same condition independently to each pair $(\mathcal{R},\mathcal{O}_{i})$. Additionally, non-convex polygons can be represented as a union of multiple convex polygons.
@@ IV-A Optimization-Based SDF Formulation in $\mathcal{M}_{\mathcal{D}}$-space
where $\|\cdot\|_{2}$ is the Euclidean norm and $(\mathbf{x}^{*},\mathbf{y}^{*})$ is the optimal solution pair to (9). The distance between sets can be reformulated as the distance between a point and a set [39] using $\text{dist}(\mathcal{R},\mathcal{O})=\text{dist}(\bm{0}_{c},\mathcal{O}\oplus(-\mathcal{R})),$ where $\mathcal{O}\oplus(-\mathcal{R})$ is their CO. Thus, we can formulate the equivalent problem in $\mathcal{M}_{\mathcal{D}}$-space as
@@
where $\bm{z}\in\mathbb{R}^{d}$ is the MTV, and =$\mathcal{R}^{\prime}=\{y\in\mathbb{R}^{d}\mid A_{r}y\leq b_{r}+A_{r}\bm{z}\}$ is the translation of $\mathcal{R}$ by $\bm{z}$. However, unlike in the minimum distance case, it is unclear how to represent the separation constraint $\mathcal{R}^{\prime}\cap\mathcal{O}=\varnothing$ as a convex constraint.
@@
where $s\in\mathbb{R}$ is the minimal translational distance required to seperate two overlapping sets. The vector $A^{\mathcal{C}}_{k}$ and scalar $b^{\mathcal{C}}_{k}$, denoting the $k$-th row $A^{\mathcal{C}}$ and $b^{\mathcal{C}}$, respectively, define the bounding hyperplanes of $\mathcal{O}^{\mathcal{C}}$. It is shown in [33, 25] that $\text{pd}(\mathcal{R},\mathcal{O})=\text{depth}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}})$.
However, while (12) gives the penetration depth, we also need the direction of separation to recover the optimal solution $\bm{z}^{*}$ of (11), which cannot be solved directly. This direction can be obtained by identifying the active constraint from the dual variable of (12). Let $k_{\text{act}}$ be the set of active indices, and $A^{\mathcal{C}}_{k_{\text{act}}}$ be the matrix containing the rows corresponding to the active constraints at the optimal solution $s^{*}$. By strong duality in LP, complementary slackness holds and implies that the nonzero dual variables $\lambda_{k_{\text{act}}}$ correspond to the active constraints.
In most cases, there is only one active constraint, meaning that $A^{\mathcal{C}}_{k_{\text{act}}}$ reduces to a single row vector, denoted as $a^{\mathcal{C}}_{k_{\text{act}}}$, which determines the direction of separation. When multiple constraints are active, $A^{\mathcal{C}}_{k_{\text{act}}}$ becomes a matrix. As the origin lies equidistant from the corresponding active constraint planes, the direction of separation can be determined by any of them. While the presence of multiple active constraints is theoretically possible, it occurs only for a measure zeros set of configurations. Moreover, should such a case occur in practice, then any of the active constraints can be selected and used individually (since all active constraints essentially represent different but equidistant paths out of the obstacle).
Once the active constraint of (12) is identified, corresponding to the hyperplane =$\mathcal{H}=\{y\in\mathbb{R}^{d}\mid a^{\mathcal{C}}_{k_{\text{act}}}y=b^{\mathcal{C}}_{k_{\text{act}}}\}$, $\bm{z}^{*}$ can be computed by projecting the origin $\bm{0}_{c}$ onto $\mathcal{H}$ [39] as
@@
The signed distance (7) is computed by solving either the QP in (10) or the LP in (12), with feasibility guaranteed by the first assumption. The equivalence of computing the signed distance in both $\mathcal{W}$- and $\mathcal{M}_{\mathcal{D}}$-spaces is demonstrated in Fig. 1. The key advantages of formulating the optimization problems in $\mathcal{M}_{\mathcal{D}}$-spaceare: (1) it enables convex optimization (LP) formulation for penetration depth computation; and (2) the minimum distance and penetration depth can be viewed as companion problems that provide a generalized measure of distance. This allows us to compute derivative of the signed distance using the same differentiable optimization framework, which will be discussed in Sec. IV-C.

### v3

@@ Abstract
Safely navigating around obstacles while respecting the dynamics, control, and geometry of the underlying system is a key challenge in robotics. Control Barrier Functions (CBFs) generate safe control policies by considering system dynamics and geometry when calculating safe forward-invariant sets. Existing CBF-based methods often rely on conservative shape approximations, like spheres or ellipsoids, which have explicit and differentiable distance functions. In this paper, we propose an optimization-defined CBF that directly considers the exact Signed Distance Function (SDF) between a polytopic robot and polytopic obstacles. Inspired by the Gilbert-Johnson-Keerthi (GJK) algorithm, we formulate both (i) minimum distance and (ii) penetration depth between polytopic sets as convex optimization problems in the space of Minkowski difference operations (the $\mathcal{M}_{\mathcal{D}}$-space). Convenient geometric properties of the $\mathcal{M}_{\mathcal{D}}$-space enable the derivatives of implicit SDF between two polytopes to be computed via differentiable optimization. We demonstrate the proposed framework in three scenarios including pure translation, initialization inside an unsafe set, and multi-obstacle avoidance. These three scenarios highlight the generation of a non-conservative maneuver, a recovery after starting in collision, and the consideration of multiple obstacles via pairwise CBF constraints, respectively. A video of all simulation animations can be found at https://youtu.be/3Dh0gtDW8bE .
@@ I Introduction
Safely navigating an environment with obstacles is a fundamental challenge in robotics. Navigational solutions must respect the dynamics and control constraints of the vehicle and the geometric constraints of the vehicle and obstacles. Control Barrier Functions (CBFs) guarantee safety for control-affine systems by providing forward-invariant safe sets and combine naturally with CLFs, yielding a sequence of pointwise quadratic programs (QP) [1, 2]. QPs are computationally efficient and well-suited to real-time applications. Further advancements in safety-critical control with CBFs have introduced formulations for high relative degree constraints [3, 4, 5, 6, 7].
@@
CBFs are used for obstacle avoidance across a range of applications [8, 9, 10, 11, 12, 13, 14], yet most CBFs approximate robots and/or obstacles by smooth shapes, i.e., hyperspheres [15] and ellipsoids [16, 17]. Hyperspheres and ellipsoids are popular because they admit closed-form, differentiable robot-obstacle distance, but they are often overly conservative approximations of real-world objects; such objects are, for example, more accurately approximated by polytopes.
@@
In this paper we present a CBF-based framework that considers polytopic robots and polytopic obstacles directly. Our method computes the exact signed distance function (SDF) between two polytopes by solving a convex optimization problem in the dual space that is defined by the Minkowski difference between the two polytopes (see Fig. 1). The resulting robot-to-obstacle SDF is differentiable almost everywhere (a.e.).
Our work builds on a foundation of prior work [18, 19, 20, 21, 22, 23] that considers continuous-time CBFs for collision avoidance between polytopes/polygons. The first study of CBFs [18] uses the Minimum Distance Function (MDF), solved by duality-based convex optimization, to construct a nonsmooth CBF (NCBF). In [18, 19] the exact derivative is applied in the case of strongly convex sets (but not polytopes, which are weakly convex), while a conservative lower bound on the CBF derivative is used for polytopes. A CBF-based filtering strategy for manipulators [20] uses the SDF and a local linear approximation of the CBF gradient (approximation error considered to ensure safety). An optimization-free smooth approx-SDF CBF is proposed in [23], where conservatism is reduced via hyperparameter tuning. In [22] CBFs are constructed using differentiable optimization of the minimum scaling factor (another metric for collision detection), and polygons are approximated by strictly convex shapes. Polygons are similarly approximated by polynomial shapes using logistic regression in [21]. Our work differs from prior work [18, 19, 20, 21, 22, 23] in that we solve the SDF via convex optimization and compute its derivative using differentiable optimization without approximation by leveraging the configuration obstacle defined via the Minkowski difference. Table I provides a conceptual comparison with existing works.
The main contributions of this paper are as follows:
II Preliminaries
II-A Control Barrier Functions
Consider a nonlinear affine control system:
$\dot{\bm{x}}=f(\bm{x})+g(\bm{x})\bm{u},$ (1)
Definition 1 (Forward invariant set):
Let $C$ be the safe set, defined as the 0-superlevel set of continuously differentiable $h:\mathcal{X}\subset\mathbb{R}^{n}\rightarrow\mathbb{R}$ :
Definition 2 (CBF [2]):
Given a set $C$ as in (2), $h(\bm{x})$ is a control barrier function (CBF) for system (1) if there exists an extended class $\mathcal{K}_{\infty}$ function [^2] An extended class $\mathcal{K}_{\infty}$ function is a function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ that is strictly increasing and $\alpha(0)=0$, see, e.g., [1]. $\alpha$ such that, for all $\bm{x}\in\mathcal{X}$,
where $L_{f}(\cdot),L_{g}(\cdot)$ denote the Lie derivatives along $f$ and $g$, respectively. It is assumed that $L_{g}h(\bm{x})\neq 0$ when $h(\bm{x})=0$.
Theorem 1 (Safety Guarantee from [2]).
Given a CBF $h$ with the associated set $C$ from (2), any Lipschitz continuous controller $\bm{u}(t)\in\mathcal{U},t\geq t_{0}$ that satisfies (3) renders the set $C$ forward invariant for (1).
Remark 1.
Our SDF is continuously differentiable w.r.t. spatial variables but (only) differentiable a.e. w.r.t. orientation (due to active-set changes) and (only) when the robot and obstacle are not in collision. Similarly, in the cases of contact/penetration, the SDF is locally Lipschitz, and thus (only) differentiable a.e., such that non-differentiability occurs on a measure-zero set (vertices, interior medial axis, rotation-induced active-set switches). The classical CBF in Def. 2 does not apply at those (measure-zero set of) points. Consequently, the proofs in this paper (which build from the classical approach) fully apply to the non-collision case for single-integrator dynamics. We conjecture that a parallel analysis can be obtained a.e. for the collision and orientation-dependent cases; however, a rigorous nonsmooth analysis such as in [31, 19] is planned for future work.
Definition 3 (Control Lyapunov function [32]):
@@ II-A Control Barrier Functions
Many existing works [2, 33, 4], unify CBF and CLF within a QP framework by relaxing the CLF with a slack variable penalized in the cost function to improve feasibility.
@@ II-B Collision Detection via Minkowski Operation
The workspace $\mathcal{W}$ of a robot is the physical (2D or 3D Euclidean) environment in which it operates, consisting of the space where the robot can move and the obstacles it must avoid. The configuration space ($\mathcal{C}$-space) represents all possible positions, orientations, etc, with each point defining a unique configuration. In contrast, the $\mathcal{M}_{\mathcal{D}}$-space is a dual of $\mathcal{W}$, and is a Euclidean space with the same (2D or 3D) dimensionality as $\mathcal{W}$. In $\mathcal{M}_{\mathcal{D}}$-space, the robot is shrunk to a point at the origin, and the obstacles are transformed to encode the robot’s geometry and configuration (see Fig. 1).
@@
where $\mathcal{R}=\{-\bm{p}\mid\bm{p}\in\mathcal{R}\}$ is the reflection of $\mathcal{R}$ through the origin, where $\bm{p}$ is a position vector of $\mathcal{R}$. Note that we refer to the Minkowski sum of $\mathcal{O}$ and -$\mathcal{R}$ as the Minkowski difference, following collision-detection convention. [^3] In fields like mathematics and image processing, the Minkowski difference has a different (alternative) definition, where $\mathcal{A}\ominus\mathcal{B}$ is the set that, when added to $\mathcal{B}$, recovers $\mathcal{A}$.
@@
While the Minkowski sum (5) and CO computation (6) are general operations that apply to any shape, it is worth emphasizing that Lem. 1 holds only for convex sets. For well-behaved non-convex sets (e.g., non-fractal) that can be decomposed into convex subsets, a CO is computed for each convex-robot-convex-obstacle pair, and Lem. 1 applied iteratively. In this paper, we consider nonempty convex polygons for both robot and obstacle.
@@ II-C Signed distance for Collision Avoidance
The signed distance is positive when $\mathcal{R}$ and $\mathcal{O}$ are disjoint and negative when they intersect. Thus, collision avoidance can be enforced by ensuring $\text{sd}(\mathcal{R},\mathcal{O})>0$. The signed distance between two sets can equivalently be expressed as the signed distance between the origin $\bm{0}_{c}$ and their CO. The closest related work [18] constructs CBFs based on implicit MDFs by posing a convex program with dual variables, and enforces the resulting NCBF with a linear program (LP) that imposes a lower bound on the CBF derivative. Moreover, [18] considers only minimum distance and not penetration depth (cases where the system starts in collision). In contrast, we propose a new framework for SDF-based CBFs that incorporates penetration depth, as detailed in Sec. IV.
@@ III Problem Formulation
The robot’s state is denoted $\bm{x}$. Both robot and obstacles are defined using the halfspace-representation (H-rep), $\mathcal{R}(\bm{x}):=\{y\in\mathbb{R}^{d}\mid A_{r}(\bm{x})y\leq b_{r}(\bm{x})\}$ and $\mathcal{O}_{i}:=\{y\in\mathbb{R}^{d}\mid A_{o_{i}}y\leq b_{o_{i}}\}$, with $A_{r}\in\mathbb{R}^{\ell_{r}\times d},\leavevmode\nobreak\ b_{r}\in\mathbb{R}^{\ell_{r}}$, $A_{o_{i}}\in\mathbb{R}^{\ell_{o_{i}}\times d},\leavevmode\nobreak\ b_{o_{i}}\in\mathbb{R}^{\ell_{o_{i}}}$, and $d\in\{2,3\}$ for 2D and 3D spaces. $\ell_{r}$ and $\ell_{o_{i}}$ denote the number of edges of the robot and the $i$-th obstacle, respectively. The corresponding COs are denoted as $\mathcal{O}_{i}^{\mathcal{C}}(\bm{x})=\{y\in\mathbb{R}^{d}\mid A^{\mathcal{C}}_{i}(\bm{x})y\leq b^{\mathcal{C}}_{i}(\bm{x})\}$, omitting the subscript $\mathcal{R}$ for brevity, where $A_{i}^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}_{i}}\times d},\leavevmode\nobreak\ b_{i}^{\mathcal{C}}\in\mathbb{R}^{\ell_{\mathcal{C}_{i}}}$, and $\ell_{\mathcal{C}_{i}}$ denotes the number of edges of the $i$-th CO.
@@
  2. 2.
    $A^{\mathcal{C}}_{i}(\bm{x}),b^{\mathcal{C}}_{i}(\bm{x})$ are differentiable a.e., and the inequality constraints provide a minimal representation of the CO. For almost all $\bm{x}\in\mathcal{X}$, $\ell_{\mathcal{C}_{i}}=\ell_{r}+\ell_{o_{i}}$. Note that, if $\mathcal{R}$ and $\mathcal{O}_{i}$ have parallel edges, the associated constraints $A^{\mathcal{C}}_{i},\leavevmode\nobreak\ b^{\mathcal{C}}_{i}$ may lose one/more rows, depending on the number of parallel edges. Yet, such degeneracy occurs only on a measure zero set and thus does not affect the optimization for almost all $\bm{x}$ (as reported in [38]).
  3. 3.
    All obstacles are static, uniformly bounded, and have nonempty interior. Also, the information about obstacles is accessible to the robot.
@@
    achieves objectives: minimizes energy consumption $J=\int_{0}^{T}\bm{u}^{\top}\bm{u}dt$, and ensures the robot reaches the desired position $(x_{d},y_{d})$,
@@
This formulation extends to multiple obstacles by applying the same safety constraint independently to each pair $(\mathcal{R},\mathcal{O}_{i})$. Additionally, non-convex polygons can be represented as a union of multiple convex polygons.
@@ IV-A Optimization-Based SDF Formulation in $\mathcal{M}_{\mathcal{D}}$-space
where $\|\cdot\|_{2}$ is the Euclidean norm and $(\mathbf{x}^{*},\mathbf{y}^{*})$ is the optimal solution pair to (9). The distance between sets can be reformulated as the distance between a point and a set [40], $\text{dist}(\mathcal{R},\mathcal{O})=\text{dist}(\bm{0}_{c},\mathcal{O}\oplus(-\mathcal{R})),$ where $\mathcal{O}\oplus(-\mathcal{R})$ is their CO. We formulate the equivalent problem in $\mathcal{M}_{\mathcal{D}}$-space as
@@
where $\bm{z}\in\mathbb{R}^{d}$ is the MTV, and $\mathcal{R}^{\prime}=\{y\in\mathbb{R}^{d}\mid A_{r}y\leq b_{r}+A_{r}\bm{z}\}$ is the translation of $\mathcal{R}$ by $\bm{z}$. However, unlike in the minimum distance case, it is unclear how to represent the separation constraint $\mathcal{R}^{\prime}\cap\mathcal{O}=\varnothing$ as a convex constraint.
@@
where $s\in\mathbb{R}$ is the minimal translational distance required to separate two overlapping sets. The vector $A^{\mathcal{C}}_{k}$ and scalar $b^{\mathcal{C}}_{k}$, denoting the $k$-th row $A^{\mathcal{C}}$ and $b^{\mathcal{C}}$, respectively, define the bounding hyperplanes of $\mathcal{O}^{\mathcal{C}}$. It is shown in [34, 25] that $\text{pd}(\mathcal{R},\mathcal{O})=\text{depth}(\bm{0}_{c},\mathcal{O}^{\mathcal{C}})$.
However, while (12) gives the penetration depth (scalar), we also need the direction of separation to recover the optimal solution $\bm{z}^{*}$ of (11), which cannot be solved directly. This direction can be obtained by identifying the active constraint from the dual variable of (12). Let $k_{\text{act}}$ be the set of active indices, and $A^{\mathcal{C}}_{k_{\text{act}}}$ be the matrix containing the rows corresponding to the active constraints at the optimal solution $s^{*}$. By strong duality in LP, complementary slackness holds and implies that nonzero dual variables $\lambda_{k_{\text{act}}}$ correspond to the active constraints.
In most cases, a single constraint is active, so $A^{\mathcal{C}}_{k_{\text{act}}}$ reduces to one row vector, denoted as $a^{\mathcal{C}}_{k_{\text{act}}}$, which uniquely determines the separation direction. When multiple constraints are active, $A^{\mathcal{C}}_{k_{\text{act}}}$ becomes a matrix. As the origin is equidistant from multiple active planes, the separating direction lies in the convex cone of their normals. The presence of multiple active constraints occurs only for a measure-zero set of configurations. Moreover, if such a case occurs in practice, any active constraint yields a valid separating direction. Once the active constraint of (12) is identified, corresponding to the hyperplane $\mathcal{H}=\{y\in\mathbb{R}^{d}\mid a^{\mathcal{C}}_{k_{\text{act}}}y=b^{\mathcal{C}}_{k_{\text{act}}}\}$, $\bm{z}^{*}$ can be computed by projecting the origin $\bm{0}_{c}$ onto $\mathcal{H}$ [40] as
@@
The signed distance (7) is computed by solving either the QP in (10) or the LP in (12), with feasibility guaranteed by the assumption 1. and Lem. 2. The equivalence of computing the signed distance in both $\mathcal{W}$- and $\mathcal{M}_{\mathcal{D}}$-spaces is shown in Fig. 1. The key advantages of formulating the optimization problems in $\mathcal{M}_{\mathcal{D}}$-space are: (1) it enables convex optimization (LP) formulation for penetration depth computation; and (2) the minimum distance and penetration depth can be viewed as companion problems that provide a generalized measure of distance. This allows us to compute derivative of the signed distance using the same differentiable optimization framework, see Sec. IV-C.